a quantity of low density gas in rigid container is initially at room temperature ( 20 degrees C ) and a particular pressure P1. If the gas Is heated to a temperature of 60 degrees C, by what factor does the pressure change.




1
Expert's answer
2021-11-22T16:15:43-0500

The volume is constant, by the combined gas law, we have:


"\frac{P}{T}=\frac{P_1}{T_1},\\\space\\
\frac{P_1}{P}=\frac{T_1}{T}=\frac{273+60}{273+20}=1.14."